OpenStudy (anonymous):

A 67.0 kg crate is being raised by means of a rope. Its upward acceleration is 3.50 m/s2. What is the force exerted by the rope on the crate? A 67.0 kg crate is being raised by means of a rope. Its upward acceleration is 3.50 m/s2. What is the force exerted by the rope on the crate? @Mathematics

OpenStudy (stormfire1):

F = M*A so just plug in M*A and solve

OpenStudy (anonymous):

incorrect!!!

OpenStudy (stormfire1):

Yea, I left out gravity...one sec

OpenStudy (anonymous):

ok

OpenStudy (stormfire1):

T = (mg) + (ma) = 67(9.8) + 67(3.5) = 891.1N

OpenStudy (anonymous):

correct!!!

OpenStudy (anonymous):

when do we normally use this formula? cos i know the normal formula to be f=ma ?

OpenStudy (stormfire1):

good...I shouldn't have made that first mistake to begin with :)

OpenStudy (stormfire1):

F=MA is used all the time. For tension, you have to add the downward force PLUS the upward force together to get the *total* force on the rope.

OpenStudy (stormfire1):

Most of the time when someone misses a physics question like this, it's because they didn't account for all the forces involved. That's why force diagrams are so important.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

A 40.0 kg crate is being lowered by means of a rope. Its downward acceleration is 3.80 m/s2. What is the force exerted by the rope on the crate?

OpenStudy (stormfire1):

Same thing really except that since the acceleration is in the same direction as the force of gravity, you subtract: T = m(g-a) = 40(9.8 - 3.8) = 240N